OpenStudy (mathhelp346):

The perimeter of a square is at least 15 cm more than the length of a side. Express the statement as an inequality, and solve for the possible length of the side.

OpenStudy (mathhelp346):

?

OpenStudy (mertsj):

p>=s+15

OpenStudy (anonymous):

\[4s \ge s+15\]

OpenStudy (mertsj):

s<=p-15

OpenStudy (mathhelp346):

so it wold be p> or = to s+15?

OpenStudy (mertsj):

Yes.

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):

\[3s \ge 15\] so \[s \ge 5\]

OpenStudy (mertsj):

Now remember that p = 4s so you could substitute 4s in place of p and have a simpler sentence because it would only have 1 variable.

OpenStudy (mathhelp346):

but what would be possible lengths?

OpenStudy (mertsj):

just like creative has shown you.

OpenStudy (mathhelp346):

ok

OpenStudy (mertsj):

Do you get it?

OpenStudy (mathhelp346):

yes

OpenStudy (mertsj):

Excellent.

OpenStudy (mathhelp346):

thanks

OpenStudy (mertsj):

yw